OpenStudy (anonymous):

Five years ago, the town of Mayberry had a population 1532. The town has grown at 11.05% per year. What is the current population?

OpenStudy (radar):

Treat it liker a deposit of $1532 with an interest rate of 11.05% compounded annually for 5 years.

OpenStudy (wolf1728):

Where the rate is .1105 Population = 1,532 * (1+r)^n (number of years) Population = 1,532 * (1.1105)^5 Population = 1,532 * 1.688856752 Population = 2,587.33